# Sufficient conditions for QMC analysis of finite elements for

### v1

@@ 3.3 Unbounded and Gevrey regular model
Although $U_{\boldsymbol{b}}$ is not a countable product of subsets of $\mathbb{R}$, it is still $\boldsymbol{\mu}_{G}$ measurable and of full Gaussian measure, i.e., $\boldsymbol{\mu}_{G}(U_{\boldsymbol{b}})=1$, see [40, Lem. 2.28]. In this setting, there holds
$\left\|\frac{\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}=\boldsymbol{b}^{\boldsymbol{\nu}},$
showing that our theory covers this framework.
4 Finite element methods that fulfill (A1)(A3)
Let us consider a mesh ${\mathcal{E}_{h}}\!=\!\{E_{1},E_{2},\dots,$
$E_{N}\}$. The union of the element boundaries is denoted by $\partial{\mathcal{E}_{h}}=\bigcup_{E\in{\mathcal{E}_{h}}}\partial E$. For functions $u_{h},v_{h}\in\prod_{E\in{\mathcal{E}_{h}}}L^{2}(E),\boldsymbol{q}_{h},\boldsymbol{r}_{h}\in\prod_{E\in{\mathcal{E}_{h}}}L^{2}(E)^{d}$, and $m_{h},\mu_{h}\in\prod_{E\in{\mathcal{E}_{h}}}L^{2}(\partial E)$, define scalar products
@@ 4 Finite element methods that fulfill (A1)(A3)
with induced norms $\|u_{h}\|_{{\mathcal{E}_{h}}}$, $\|m_{h}\|_{\partial{\mathcal{E}_{h}}}$, and $\|\boldsymbol{q}_{h}\|_{{\mathcal{E}_{h}}}$. In this section, we show that several (classes of) finite element methods verify Assumptions (A1), (A2) and (A3).
@@ 4.1 Mixed methods (MMs)
We consider mixed methods (MM) where $u_{h}(\boldsymbol{y})$ lies in a finite-dimensional subspace $V_{h}$ of $L^{2}(D)$, the flux approximation $\boldsymbol{q}_{h}(\boldsymbol{y})$ lies in a finite-dimensional subspace $Q_{h}$ of $H(\operatorname{div},D)$ equipped with norm
(21) $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}=\|\sqrt{a(\boldsymbol{y})}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{{\mathcal{E}_{h}}}.$
The spaces $V_{h}$ and $Q_{h}$ for MM are selected such that there is $\beta>0$ with
(22) $\sup_{\boldsymbol{r}_{h}\in Q_{h}}\frac{(\nabla\cdot\boldsymbol{r}_{h},v_{h})_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}\geq\beta\|v_{h}\|_{D}\qquad\text{ holds for all }v_{h}\in V_{h}.$
@@
(23a) $(a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{D}-(u_{h}(\boldsymbol{y}),\nabla\cdot\boldsymbol{r}_{h})_{D}$ $=G(\boldsymbol{r}_{h})$ $\text{ for all }\boldsymbol{r}_{h}\in Q_{h},$
@@
and $F(v_{h})=(f,v_{h})_{D}$, $G(\boldsymbol{r}_{h})=0$. Examples of methods in which the conditions on $V_{h}$ and $Q_{h}$ are satisfied include the Raviart–Thomas (RT) method on simplices or $d$-rectangles and Brezzi–Douglas–Marini (BDM) method on simplices. The main result of this subsection is as follows.
@@
For any $M$ satisfying the inf-sup condition (22), Assumptions (A1), (A2) and (A3) hold.
@@
It is obvious from (21) that Assumption (A1) holds with $C_{E}=1$. By Lemma 4.3 below, we see that Assumption (A2) holds with $C_{S}=1/\beta$. By Lemma 4.5, Assumption (A3) holds with $C_{R}=1$.
@@
The system (23) has a unique solution for any $\boldsymbol{y}$ and any $G$ and $F$ in the dual space of $Q_{h}$ and $V_{h}$, respectively. When $G=0$ and $F(v_{h})=(f,v_{h})_{D}$, the solution $(u_{h}(\boldsymbol{y}),\boldsymbol{q}_{h}(\boldsymbol{y}))\in V_{h}\times Q_{h}$ satisfies
$\frac{\beta\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}}{\sqrt{\overline{a}(\boldsymbol{y})}}\leq\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq\frac{\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{\mathcal{E}_{h}}}{\beta}.$
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{Q_{h}}=F(u_{h}(\boldsymbol{y}))\leq\|f\|_{\mathcal{E}_{h}}\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}.$
@@
$\beta\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}$ $\leq\sup_{\boldsymbol{r}_{h}\in Q_{h}}\frac{(\nabla\cdot\boldsymbol{r}_{h},u_{h}(\boldsymbol{y}))_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}=\sup_{\boldsymbol{r}_{h}\in Q_{h}}\frac{(a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}$
$\leq\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\sup_{\boldsymbol{r}_{h}\in Q_{h}}\frac{\|\boldsymbol{r}_{h}\|_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}\leq\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}},$
@@
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}\|_{Q_{h}}.$
@@
(24a) $(a(\boldsymbol{y})\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{D}-(\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}u_{h}(\boldsymbol{y}),\nabla\cdot\boldsymbol{r}_{h})_{D}$ $=G(\boldsymbol{r}_{h})$ $\text{ for all }\boldsymbol{r}_{h}\in Q_{h},$
@@
(25b) $\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}\|_{Q_{h}}\|\boldsymbol{r}_{h}\|_{Q_{h}}.$
@@ 4.2 Hybridizable and embedded discontinuous Galerkin (HDG and EDG)
The hybridizable discontinuous Galerkin (HDG) method identifies $u_{h}(\boldsymbol{y})\in V_{h}\subset H^{1}({\mathcal{E}_{h}})$, its flux $\boldsymbol{q}_{h}(\boldsymbol{y})\in Q_{h}\subset H(\operatorname{div},{\mathcal{E}_{h}})^{d}$ with norm
(26) $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}=\|\sqrt{a(\boldsymbol{y})}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{{\mathcal{E}_{h}}}+\|\sqrt{\tau}(u_{h}-m_{h})\|_{\partial{\mathcal{E}_{h}}},$
and $m_{h}\in M_{h}\subset L^{2}(\mathcal{F}_{h})$ with $m_{h}|_{\partial D}=0$ such that
(27a) $(a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{\mathcal{E}_{h}}-(u_{h}(\boldsymbol{y}),\nabla\cdot\boldsymbol{r}_{h})_{\mathcal{E}_{h}}+\langle m_{h}(\boldsymbol{y}),\boldsymbol{r}_{h}\cdot\boldsymbol{n}\rangle_{\partial{\mathcal{E}_{h}}}$ $=G(\boldsymbol{r}_{h})$ $\text{for all }\boldsymbol{r}_{h}\in Q_{h},$
@@
where we choose the spaces $V_{h}$ and $Q_{h}$, such that there is $\beta>0$ with
(28) $\sup_{\boldsymbol{r}_{h}\in Q_{h}}\frac{(\nabla\cdot\boldsymbol{r}_{h},v_{h})_{E}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},E)}}\geq\beta\|v_{h}\|_{E}\qquad\text{ holds for all }v_{h}\in V_{h}\text{ and all }E\in{\mathcal{E}_{h}},$
and the parameter $\tau=0$. Alternatively, we choose the parameter $\tau>0$ and can freely choose the spaces $V_{h}$ and $Q_{h}$. The former setting includes the hybridized Raviart–Thomas (RT-H) method (on simplices and $d$-rectangles) and the hybridized Brezzi–Douglas–Marini (BDM-H) method (on simplices). At the same time, the latter approach covers the hybridizable local discontinuous Galerkin scheme (LDG-H) and the class of embedded discontinuous Galerkin (EDG) schemes. Again, $F(v_{h})=(f,v_{h})_{D}$, $G(\boldsymbol{r}_{h})=0$.
Notably, (27) can be statically condensed to form a symmetric positive definite system of linear equations; see [5, Thm. 2.1], and we have the following relation to the mixed method:
@@
The RT-H and BDM-H methods produce the same approximations as the RT and BDM methods with $m_{h}=u_{h}$ on the mesh faces, respectively. That is, assumptions (A1), (A2), and (A3) hold.
@@
The inf-sup condition (28) allows us to bound $\|u_{h}(\boldsymbol{y})\|\lesssim\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}$, which is needed to show show (A2), see the proof of Lemma 4.3. For HDG methods, we assume a similar inequality, i.e.,
(29) $\|u_{h}\|_{\mathcal{E}_{h}}\leq C_{B}\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}\|_{Q_{h}},$
which can (in general) be established from (27a) and (28), or from a generalization of Poincaré’s inequality, or investigating the eigenvalues of HDG/EDG methods. If we use the BDM-H, RT-H, LDG-H, or EDG methods on simplicial or quadrilateral meshes, (29) is a direct consequence of [33, Thm. B.3] or (less-direct) of [1]. Relation (29) is also vital in the theory of multigrid methods for HDG schemes, see [34, (LS2) & (LS6), shown in Sect. 6]. Moreover, [1] allows to extend the result for LDG-H and EDG to certain non-simplicial/non-quadrilateral meshes.
@@
For any HDG method, Assumptions (A1) and (A3) hold. If additionally, (29) holds, then also (A2) holds.
@@
The validity of (A1) follows directly from (26) with $C_{E}=1$, and (A3) is shown with $C_{R}=1$ in Lemma 4.13. Finally, (A2) is demonstrated in Lemma 4.11 with $C_{S}=C_{B}$.
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{\mathcal{E}_{h}}.$
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{Q_{h}}=F(u_{h}(\boldsymbol{y}))\leq\|f\|_{\mathcal{E}_{h}}\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}\lesssim\|f\|_{\mathcal{E}_{h}}\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}$
@@
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{Q_{h}}.$
@@
The above considerations can also be extended to hybrid-high-order methods, for which the stabilizing boundary integral (the one involving $\tau$) is replaced by an abstract (stabilizing) bilinear form and (29) follows from [9, (HM1), (HM3) or (HM6)].

### v2

@@ 3.3 Unbounded and Gevrey regular model
Although $U_{\boldsymbol{b}}$ is not a countable product of subsets of $\mathbb{R}$, it is still $\boldsymbol{\mu}_{G}$ measurable and of full Gaussian measure, i.e., $\boldsymbol{\mu}_{G}(U_{\boldsymbol{b}})=1$, see [SchwabG11, Lem. 2.28]. In this setting, there holds
$\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{y})=a(\boldsymbol{y})\prod_{j\in{\rm supp}(\boldsymbol{\nu})}\psi_{j}(\boldsymbol{x})^{\nu_{j}}$}
and hence
$\left\|\frac{\partial_{\boldsymbol{y}}^{\boldsymbol{\nu}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}{{\leq}\,}\boldsymbol{b}^{\boldsymbol{\nu}},$
showing that the lognormal setting satisfies the ratio bound in Assumption 5.
4 Finite element methods that fulfill Assumptions 13
The purpose of this section is to demonstrate how the abstract assumptions introduced in Section 3.1 can be verified for concrete finite element discretizations of the balance law. Rather than treating these assumptions as purely formal conditions, it is useful to interpret them operationally in terms of familiar ingredients from finite element analysis. 1 expresses that the discrete norm $\|\cdot\|_{\mathcal{W}_{h}}$ controls the physical energy of the flux, i.e., that the chosen method provides an energy norm dominating the $L^{2}(D)$ norm of the flux weighted by the coefficient. In practice this follows from standard boundedness properties of the bilinear forms defining the method and from the definition of the discrete norm. 2 represents the stability of the discretization: the discrete flux must remain bounded in the $\mathcal{W}_{h}$-norm in terms of the data and the coefficient. For mixed and hybridizable methods this typically follows from a discrete inf–sup condition together with boundedness of the coefficient field. 3 is the key parametric derivative recursion required for the QMC analysis. Operationally, it is obtained by differentiating the discrete equations with respect to the parameter variables and using stability of the discrete problem to bound the resulting terms; the constants appearing in the recursion correspond directly to the stability constant of the method and the bounds on derivatives of the coefficient in 5
With this interpretation in mind, verifying the sufficient conditions reduces to three standard steps for each numerical scheme: (i) identify the discrete flux space $\mathcal{W}_{h}$ and norm $\|\cdot\|_{\mathcal{W}_{h}}$ so that 1 holds; (ii) establish stability of the discrete formulation, typically via an inf–sup argument or coercivity estimate yielding 2; and (iii) differentiate the discrete equations with respect to the parameters and apply the stability estimate to obtain the recursive bound 3 The remaining assumptions involve either the quantity of interest or the coefficient model and are therefore largely method-independent. The derivations below illustrate these steps for several representative discretizations—including conforming finite elements, mixed methods, and hybridizable discontinuous Galerkin schemes—while explicitly identifying how their stability constants and norms enter the abstract framework.
Let us consider a mesh ${\mathcal{E}_{h}}\!=\!\{E_{1},E_{2},\dots,E_{N}\}$. The union of the element boundaries is denoted by $\partial{\mathcal{E}_{h}}=\bigcup_{E\in{\mathcal{E}_{h}}}\partial E$. For functions $u_{h},v_{h}\in\prod_{E\in{\mathcal{E}_{h}}}L^{2}(E),\boldsymbol{q}_{h},\boldsymbol{r}_{h}\in\prod_{E\in{\mathcal{E}_{h}}}L^{2}(E)^{d}$, and $m_{h},\mu_{h}\in\prod_{E\in{\mathcal{E}_{h}}}L^{2}(\partial E)$, define scalar products
@@ 4 Finite element methods that fulfill Assumptions 13
with induced norms $\|u_{h}\|_{{\mathcal{E}_{h}}}$, $\|m_{h}\|_{\partial{\mathcal{E}_{h}}}$, and $\|\boldsymbol{q}_{h}\|_{{\mathcal{E}_{h}}}$. In this section, we show that several (classes of) finite element methods verify 1, 2 and 3.
@@ 4.1 Mixed methods (MMs)
We consider mixed methods (MM) where $u_{h}(\boldsymbol{y})$ lies in a finite-dimensional subspace $V_{h}$ of $L^{2}(D)$, the flux approximation $\boldsymbol{q}_{h}(\boldsymbol{y})$ lies in a finite-dimensional subspace $\mathcal{W}_{h}$ of $H(\operatorname{div},D)$ equipped with norm
(21) $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}=\|\sqrt{a(\boldsymbol{y})}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{{\mathcal{E}_{h}}}.$
The spaces $V_{h}$ and $\mathcal{W}_{h}$ for MM are selected such that there is $\beta>0$ with
(22) $\sup_{\boldsymbol{r}_{h}\in\mathcal{W}_{h}}\frac{(\nabla\cdot\boldsymbol{r}_{h},v_{h})_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}\geq\beta\|v_{h}\|_{D}\qquad\text{ holds for all }v_{h}\in V_{h}.$
@@
(23a) $(a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{D}-(u_{h}(\boldsymbol{y}),\nabla\cdot\boldsymbol{r}_{h})_{D}$ $=G(\boldsymbol{r}_{h})$ $\text{ for all }\boldsymbol{r}_{h}\in\mathcal{W}_{h},$
@@
and $F(v_{h})=(f,v_{h})_{D}$, $G(\boldsymbol{r}_{h})=0$. Examples of methods in which the conditions on $V_{h}$ and $\mathcal{W}_{h}$ are satisfied include the Raviart–Thomas (RT) method on simplices or $d$-rectangles and Brezzi–Douglas–Marini (BDM) method on simplices. The main result of this subsection is as follows.
@@
For any $M$ satisfying the inf-sup condition (22), 1, 2 and 3 hold.
@@
It is obvious from (21) that 1 holds with $C_{E}=1$. By Lemma 4.3 below, we see that 2 holds with $C_{S}=1/\beta$. By Lemma 4.5, 3 holds with $C_{R}=1$.
@@
The system (23) has a unique solution for any $\boldsymbol{y}$ and any $G$ and $F$ in the dual space of $\mathcal{W}_{h}$ and $V_{h}$, respectively. When $G=0$ and $F(v_{h})=(f,v_{h})_{D}$, the solution $(u_{h}(\boldsymbol{y}),\boldsymbol{q}_{h}(\boldsymbol{y}))\in V_{h}\times\mathcal{W}_{h}$ satisfies
$\frac{\beta\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}}{\sqrt{\overline{a}(\boldsymbol{y})}}\leq\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq\frac{\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{\mathcal{E}_{h}}}{\beta}.$
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{\mathcal{W}_{h}}=F(u_{h}(\boldsymbol{y}))\leq\|f\|_{\mathcal{E}_{h}}\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}.$
@@
$\beta\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}$ $\leq\sup_{\boldsymbol{r}_{h}\in\mathcal{W}_{h}}\frac{(\nabla\cdot\boldsymbol{r}_{h},u_{h}(\boldsymbol{y}))_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}=\sup_{\boldsymbol{r}_{h}\in\mathcal{W}_{h}}\frac{(a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}$
$\leq\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\sup_{\boldsymbol{r}_{h}\in\mathcal{W}_{h}}\frac{\|\boldsymbol{r}_{h}\|_{D}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},D)}}\leq\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}},$
@@
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}\|_{\mathcal{W}_{h}}.$
@@
(24a) $(a(\boldsymbol{y})\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{D}-(\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}u_{h}(\boldsymbol{y}),\nabla\cdot\boldsymbol{r}_{h})_{D}$ $=G(\boldsymbol{r}_{h})$ $\text{ for all }\boldsymbol{r}_{h}\in\mathcal{W}_{h},$
@@
(25b) $\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}\|_{\mathcal{W}_{h}}\|\boldsymbol{r}_{h}\|_{\mathcal{W}_{h}}.$
@@ 4.2 Hybridizable and embedded discontinuous Galerkin (HDG and EDG)
The hybridizable discontinuous Galerkin (HDG) method identifies $u_{h}(\boldsymbol{y})\in V_{h}\subset H^{1}({\mathcal{E}_{h}})$, its flux $\boldsymbol{q}_{h}(\boldsymbol{y})\in\mathcal{W}_{h}\subset H(\operatorname{div},{\mathcal{E}_{h}})^{d}$ with norm
(26) $\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}=\|\sqrt{a(\boldsymbol{y})}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{{\mathcal{E}_{h}}}+\|\sqrt{\tau}(u_{h}{(\boldsymbol{y})}-m_{h}{(\boldsymbol{y})})\|_{\partial{\mathcal{E}_{h}}},$
and $m_{h}{(\boldsymbol{y})}\in M_{h}\subset L^{2}(\mathcal{F}_{h})$ with $m_{h}{(\boldsymbol{y})}|_{\partial D}=0$ such that
(27a) $(a(\boldsymbol{y})\boldsymbol{q}_{h}(\boldsymbol{y}),\boldsymbol{r}_{h})_{\mathcal{E}_{h}}-(u_{h}(\boldsymbol{y}),\nabla\cdot\boldsymbol{r}_{h})_{\mathcal{E}_{h}}+\langle m_{h}(\boldsymbol{y}),\boldsymbol{r}_{h}\cdot\boldsymbol{n}\rangle_{\partial{\mathcal{E}_{h}}}$ $=G(\boldsymbol{r}_{h})$ $\text{for all }\boldsymbol{r}_{h}\in\mathcal{W}_{h},$
@@
where we choose the spaces $V_{h}$ and $\mathcal{W}_{h}$, such that there is $\beta>0$ with
(28) $\sup_{\boldsymbol{r}_{h}\in\mathcal{W}_{h}}\frac{(\nabla\cdot\boldsymbol{r}_{h},v_{h})_{E}}{\|\boldsymbol{r}_{h}\|_{H(\operatorname{div},E)}}\geq\beta\|v_{h}\|_{E}\qquad\text{ holds for all }v_{h}\in V_{h}\text{ and all }E\in{\mathcal{E}_{h}},$
and the parameter $\tau=0$. Alternatively, we choose the parameter $\tau>0$ and can freely choose the spaces $V_{h}$ and $\mathcal{W}_{h}$. The former setting includes the hybridized Raviart–Thomas (RT-H) method (on simplices and $d$-rectangles) and the hybridized Brezzi–Douglas–Marini (BDM-H) method (on simplices). At the same time, the latter approach covers the hybridizable local discontinuous Galerkin scheme (LDG-H) and the class of embedded discontinuous Galerkin (EDG) schemes. Again, $F(v_{h})=(f,v_{h})_{D}$, $G(\boldsymbol{r}_{h})=0$.
Notably, (27) can be statically condensed to form a symmetric positive definite system of linear equations; see [CockburnGL09, Thm. 2.1], and we have the following relation to the mixed method:
@@
The RT-H and BDM-H methods produce the same approximations as the RT and BDM methods with $m_{h}=u_{h}$ on the mesh faces, respectively. That is, 1, 2, and 3 hold.
@@
The inf-sup condition (28) allows us to bound $\|u_{h}(\boldsymbol{y})\|\lesssim\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}$, which is needed to show show 2, see the proof of Lemma 4.3. For HDG methods, we assume a similar inequality, i.e.,
(29) $\|u_{h}\|_{\mathcal{E}_{h}}\leq C_{B}\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}\|_{\mathcal{W}_{h}},$
which can (in general) be established from (27a) and (28), or from a generalization of Poincaré’s inequality, or investigating the eigenvalues of HDG/EDG methods. If we use the BDM-H, RT-H, LDG-H, or EDG methods on simplicial or quadrilateral meshes, (29) is a direct consequence of [LuMR23, Thm. B.3] or (less-direct) of [Brenner03]. Relation (29) is also vital in the theory of multigrid methods for HDG schemes, see [LuRK22a, (LS2) & (LS6), shown in Sect. 6]. Moreover, [Brenner03] allows to extend the result for LDG-H and EDG to certain non-simplicial/non-quadrilateral meshes.
@@
For any HDG method, 1 and 3 hold. If additionally, (29) holds, then also 2 holds.
@@
The validity of 1 follows directly from (26) with $C_{E}=1$, and 3 is shown with $C_{R}=1$ in Lemma 4.13. Finally, 2 is demonstrated in Lemma 4.11 with $C_{S}=C_{B}$.
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq C_{S}\sqrt{\overline{a}(\boldsymbol{y})}\|f\|_{\mathcal{E}_{h}}.$
@@
$\|\boldsymbol{q}_{h}(\boldsymbol{y})\|^{2}_{\mathcal{W}_{h}}=F(u_{h}(\boldsymbol{y}))\leq\|f\|_{\mathcal{E}_{h}}\|u_{h}(\boldsymbol{y})\|_{\mathcal{E}_{h}}\lesssim\|f\|_{\mathcal{E}_{h}}\sqrt{\overline{a}(\boldsymbol{y})}\|\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}$
@@
$\|\partial^{\boldsymbol{\nu}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}\leq\sum_{\boldsymbol{0}\neq\boldsymbol{m}\leq\boldsymbol{\nu}}\begin{pmatrix}\boldsymbol{\nu}\\ \boldsymbol{m}\end{pmatrix}\left\|\frac{\partial^{\boldsymbol{m}}_{\boldsymbol{y}}a(\boldsymbol{y})}{a(\boldsymbol{y})}\right\|_{L^{\infty}(D)}\|\partial^{\boldsymbol{\nu}-\boldsymbol{m}}_{\boldsymbol{y}}\boldsymbol{q}_{h}(\boldsymbol{y})\|_{\mathcal{W}_{h}}.$
@@
The above considerations can also be extended to hybrid-high-order methods, for which the stabilizing boundary integral (the one involving $\tau$) is replaced by an abstract (stabilizing) bilinear form and (29) follows from [DiPietroDKMR24, (HM1), (HM3) or (HM6)].